# DefendCLI: {Command-Line} Driven Attack Provenance Examination

### v1

@@ IV-B2 Attack Irrelevance Reduction
Figure 7: Interphase attack association: a Leiden community detection method is used to correlate interphase attacks.
IV-B3 Attack-Chain Causal Inference
Identifying deviations that align with plausible attack inferences remains one of the most challenging problems in attack provenance analysis. To address this, we adopt explainable and practical techniques to isolate abnormal instances embedded within the inference process. Figure 6 illustrates our approach for identifying critical elements within the complete provenance graph, enabling focused and interpretable anomaly detection.
$\blacksquare$ How can we identify the most unusual elements traversed through reachable paths? Attack-relevant elements are often rare and obscured by a large volume of benign activities. To surface these critical nodes, we leverage the PageRank algorithm [31], originally developed to rank web pages by assessing their relative importance based on link structures. In this context, PageRank models the provenance graph as a network, where an edge from node $A$ to node $B$ is interpreted as a “vote” indicating the significance of $B$. Nodes with fewer incoming connections are considered more anomalous, reflecting their rarity within the observed activity flow.
Mathematically, the PageRank ($PR$) of a node is computed by summing the $PR$ values of all nodes linking to it, each divided by the number of outgoing links from the source node, and then scaled by a damping factor. This process is defined as:
$PR(v)=(1-d)+d\sum_{u\in\text{In}(v)}\frac{PR(u)}{L(u)},$
where $d$ is the damping factor (typically set to 0.85), $\text{In}(v)$ is the set of nodes linking to node $v$, and $L(u)$ is the number of outgoing links from node $u$. The algorithm operates iteratively, recalculating scores across the graph until convergence.
Unlike its traditional application, we normalize the final $PR$ scores to reflect rareness rather than prominence. This inversion highlights nodes that are less connected and potentially anomalous, enabling the effective identification of suspicious or atypical attack elements within the provenance graph.
$\blacksquare$ How can we identify the root cause of an attack among all accessible elements? [32] Identifying the root cause is essential for understanding attack progression. However, we do not define the root cause merely as the origin point in a source-to-destination path. Instead, we conceptualize it as a pivotal bridge node that connects early-stage activities to subsequent malicious behaviors.
To uncover such bridge nodes, we employ the shortest-path betweenness centrality algorithm [32], which quantifies the importance of a node based on the number of shortest paths that traverse it. Specifically, this measure captures how frequently a node serves as an intermediary on the shortest paths between all pairs of nodes in the graph. In the context of attack provenance, such nodes often correspond to compromised services or applications [33] that enable lateral movement or privilege escalation.
The computation involves the following steps: (1) Identify all node pairs $(s,t)$, where $s$ is a source and $t$ a destination. (2) For each pair, compute all shortest paths between $s$ and $t$. (3) For each path, count the occurrences of node $v$ along the path, excluding cases where $v=s$ or $v=t$. (4) Aggregate these counts across all pairs to obtain the betweenness centrality of node $v$.
This metric enables the identification of structurally significant nodes that are critical to the attack’s progression, thereby supporting effective root cause analysis in complex networked environments.
Figure 8: Explainable anomaly detection mechanism for unknown attacks: Ranked InfoPath will be used for the anomaly detection. SimHash recommender will examine the difference in command-line executions. Finally, the Police-Witness Voting Policy will select the most anomalous InfoPath for evaluation.
Figure 9: Example of SimHashed command-line executions: the red box of $\blacksquare$ is malicious.
$\blacksquare$ How can we integrate both causal inference methods for enhanced analysis? Both PageRank and shortest-path betweenness centrality evaluate the significance of nodes within the provenance graph, yielding node-level importance scores. To leverage the strengths of both methods within the detection framework, we transform node weights into edge weights, an essential step for enabling InfoPaths traversal during anomaly detection.
Given a pair of nodes $m$ and $n$ connected by an edge, we compute the corresponding Edge Weight ($EW_{(m,n)}$) as the sum of their individual node scores:
$EW_{(m,n)}=PR_{(m)}+PR_{(n)}+CB_{(m)}+CB_{(n)},$
where $PR$ denotes the PageRank score and $CB$ represents the shortest-path betweenness centrality score of a node.
To ensure comparability and stability across the graph, all edge weights are subsequently normalized to the range $[0,1]$. This normalization facilitates consistent interpretation and supports a unified, edge-centric approach to anomaly detection within the attack provenance framework.
IV-B4 Interphase Attack Association
Correlating multi-stage attacks within a single group poses a significant challenge, as it requires a holistic understanding of the full attack chain. Establishing meaningful associations between disparate phases demands both sufficient and reliable contextual information to uncover subtle dependencies across stages.
To address this challenge, we adopt a graph partitioning strategy using community detection to group strongly connected and semantically related nodes while isolating unrelated components. This is commonly achieved via the Louvain algorithm [34], a modularity optimization method frequently applied in attack graph analysis [35]. Louvain iteratively optimizes modularity until a stable partitioning is reached.
However, a well-documented limitation of Louvain [36] is its tendency to produce arbitrarily weakly connected communities. In extreme cases, the resulting communities may become disjoint after successive iterations, leading to inaccurate or incomplete data extraction, as reported in Depcomm [11]. Such fragmentation impairs the reliability of inter-phase correlation.
To mitigate this issue, DefendCLI replaces Louvain with the Leiden algorithm [37] for community detection, as illustrated in Figure 7. Empirical studies [38] have shown that Leiden not only improves partition quality by preserving community cohesion but also enhances computational efficiency, enabling faster analysis of large-scale graphs.
Furthermore, each detected community is assigned a community score, which is subsequently incorporated into the InfoPaths search process to reinforce the correlation of multi-stage attack behaviors across the provenance graph.
2
Command-line Differentiation Evaluation (Known Facts). Based on real-world feedback, it is inadvisable to disregard expert knowledge during early-stage detection.
$\blacksquare$ How can we detect known attack behaviors? To support practical deployment, we evaluate detected system process chains using Sigma rules [24] via SIEM platforms (e.g., Elasticsearch). If a matching pattern is identified, the process chain is labeled as malicious and assigned a High-Risk Impact with an associated Risk Score ($RS$).
$\blacksquare$ How can we detect known defense evasion techniques? Process chains are further analyzed using a curated set of evasion rules [25, 26]. These rules adjust the weights in the attack provenance graph based on rule confidence levels. Evasion indicators are categorized into Medium-Risk and Low-Risk classes, each assigned a corresponding $RS$.
$\blacksquare$ How can we integrate both methods into attack provenance analysis? Following command-line evaluation, identified patterns are mapped to the provenance graph using regular expressions [39]. For each matching edge between nodes $m$ and $n$, the Refined Edge Weight ($ReEW_{(m,n)}$) is computed as:
$ReEW_{(m,n)}=RS\times EW_{(m,n)},$
where $EW_{(m,n)}$ is the original edge weight. Updated edges adopt the new weight, while unmatched nodes and edges retain their original values.
IV-C Attack-Evidence Awareness
This module comprises two primary components: (1) explainable anomaly detection to identify unknown attacks and (2) automated threat report generation with actionable insights. By analyzing accumulated attack clauses, these components work in tandem to enhance understanding of attack patterns and support effective threat identification.
3
Explainable Anomaly Detection (Unknown Attacks). To detect novel attacks based on accumulated clauses, we design an explainable anomaly detection framework, as illustrated in Figure 8.
IV-C1 Information Retrieval (InfoPath Search)
A central challenge in attack provenance analysis is determining how to retrieve malicious information from large volumes of provenance data using attack clauses. To this end, we propose InfoPath Search, a retrieval method designed to efficiently extract relevant attack-related information.
$\blacksquare$ How can we retrieve malicious information based on attack clauses? InfoPaths are derived from each partitioned cluster based on interphase attack associations. We employ the Dijkstra algorithm [40] to compute the shortest paths between node pairs.
Traversal begins from nodes with zero in-degree and terminates at nodes with zero out-degree, when applicable; otherwise, all paths are considered. This approach ensures that, in a directed provenance graph, the shortest InfoPaths reflect plausible causal attack chains from source to sink.
$\blacksquare$ How does information retrieval relate to causal inference and command-line evaluation? In our framework, higher ($ReEW$) denote higher significance in causal inference. However, since shortest-path algorithms prioritize lower weights, we invert the edge weights using $\frac{1}{ReEW}$ to ensure that rare and anomalous paths are prioritized. This transformation aligns the retrieval process with the objective of uncovering unusual and potentially malicious patterns.
$\blacksquare$ How does information retrieval support interphase attack correlation? Shorter InfoPaths typically indicate more suspicious activity. To enhance interphase attack correlation, path lengths are further adjusted using the community score (CS), favoring paths that traverse high-impact communities, that is defined as $\frac{1}{ReEW\times CS}$. The final InfoPath rankings are based on two factors: (i) effective path length and (ii) the diversity of system processes involved. This accounts for the practical observations that malware often leverages various system processes to evade detection [41].
IV-C2 Command-Line Differentiation Evaluation Based on Unknown Facts (SimHash Method)
To improve the identification of novel attacks, we propose a framework that delivers an impartial anomaly detection system. This framework bridges the gap between algorithmic uncertainty and the capability to detect previously unseen command-line activities.
$\blacksquare$ How can we detect subtle context deviations to effectively address complex multi-stage attacks? Previous research often overlooks the impact of varying embedding techniques on security datasets, leading to reduced sensitivity across diverse data types. This oversight diminishes the granularity of detection and restricts context awareness to minor deviations, an essential limitation when addressing novel and sophisticated attacks.
To address this issue, we utilize the SimHash [5] algorithm to randomly assess similarity within post-embedding command-line data. Our approach incorporates embedding techniques such as Word2Vec [42], Doc2Vec [43], and FastText [44], which have been employed in previous studies. This integration aims to mitigate the impact of vectorization on different types of security audit data.
SimHash generates a unique hash value for each embedded command-line entry, as shown in Figure 9. The process begins by transforming command-line data into vector representations using Word2Vec [42], Doc2Vec [43], and FastText [44], which capture semantic relationships and contextual dependencies within command sequences.
Then, a weighted hashing function is applied to generate a fixed-length binary signature for each embedded command-line entry. This technique ensures that similar command-line sequences produce hash values with minimal Hamming distance, whereas significantly different sequences result in more distinct hashes. By computing and analyzing these distances, the approach quantifies the similarity between different command-line entries.
Finally, the embedding method that best distinguishes the differences among all evaluated command-line inputs based on SimHash similarity scores is selected for anomaly detection. This process ensures that vectorization effectively preserves critical distinctions within security audit data, enhancing the detection of novel attacks while reducing unnecessary false positives.
$\blacksquare$ How can we objectively detect anomalies based on command-line behavioral differentiation? Recent research trends favor the use of neural networks for anomaly detection. However, these methods often require large volumes of high-quality training data and substantial computing resources. In contrast, we believe that a cost-effective approach is well-suited for attack provenance examination, which typically demands lightweight on-site deployment. To achieve this goal, we developed a multifactor-based detection mechanism inspired by criminal investigation methods.
Various classifiers (IsolationForest [45], LocalOutlierFactor [46], OneClassSVM [47], EllipticEnvelope [48], KMeans [49], and GaussianMixture [50]) then respond based on their respective algorithms, ensuring diverse analytical perspectives. This approach mitigates algorithmic bias by preventing over-reliance on similar features. The final decision is determined through a binary voting system [51], where the suspect receiving the most votes is identified.
In this context, suspects are represented by InfoPaths embedded with command-line vectors, while multiple classifiers provide supporting evidence, culminating in the voting process. Ultimately, InfoPaths identified as malicious and exhibiting significant command-line differentiation are selected for further analysis.
TABLE II: Effect of Hyperparameters on Adaptive System
Hyperparameters Positive Effect Negative Effect
$C$ $\uparrow C\uparrow$ DR $\uparrow C\uparrow$ FAR
RSRS $\uparrow RS\downarrow$ FAR $\uparrow RS\downarrow$ DR
NNNN $\uparrow NN\uparrow$ DR $\uparrow NN\uparrow$ FAR
VSVS $\uparrow VS\uparrow$ DR $\uparrow VS\uparrow$ FAR
IV-C3 Adaptive Interface for Dynamic Environments
Machine learning-based detection mechanisms often demonstrate limited effectiveness in real-world environments due to the dynamic nature of software and hardware ecosystems. Frequent updates and evolving system behavior can degrade model performance over time, necessitating retraining and continuous evaluation, both of which introduce significant human and computational overhead.
To mitigate these challenges, we propose an adaptive interface that allows external users to adjust key hyperparameters that influence False Alarm Rate (FAR) and Detection Rate (DR), with particular emphasis on adjusting classification thresholds. This user-configurable approach reduces the opacity of AI-based systems and supports improved operational alignment with diverse deployment conditions.
Through extensive experiments, we identify several hyperparameters that significantly impact system performance. These are summarized in Table II, and can be tuned according to local environmental characteristics. Automated tuning, while convenient, often leads to increased false alarms in heterogeneous settings, reinforcing the need for manual control.
$\blacksquare$ Influence of Risk Score (RS) based on known attack facts: As described in Section IV-B4, expert knowledge is incorporated through command-line differentiation, categorizing risk into three levels: High (H), Medium (M), and Low (L), each associated with a corresponding RS. These default values are set as: {H = 2.5, M = 2.0, L = 1.5}, reflecting the confidence level in known attack indicators.
$\blacksquare$ Influence of Contamination ($C$) based on audit data quality: Real-world audit data often contain inconsistencies, making it difficult to accurately distinguish between benign and malicious activity. To account for data imbalance, we introduce the Contamination parameter $C$, which is integrated into supported classifiers to control for class skew. The default value $C=0.5$ represents a moderate assumption of data imbalance.
$\blacksquare$ Influence of Number of Neighbors (NN) and Vector Size (VS): Data intensity varies between environments, which requires flexibility in classifier configuration. The Number of Neighbors ($NN$) and Vector Size ($VS$) parameters help scale detection strength. Lower values are recommended for smaller datasets, while higher values are preferable in high-volume environments. The default settings are $NN=300$ and $VS=60$.
4
Report Generation and Recommendation. Alarm fatigue, a widespread problem in cybersecurity, occurs when security analysts are inundated with excessive alerts, resulting in missed critical threats and delayed responses. To mitigate this challenge, we introduce a snapshot alarm mechanism powered by a GPT model enhanced through a Retrieval-Augmented Generation (RAG) framework based on GPT-4o. The RAG system incorporates a curated set of trusted security reports [52, 53, 54], significantly improving the model’s ability to detect, contextualize, and prioritize emerging threats.
Figure 10 illustrates the GPT model recommendation workflow. The mechanism ranks all suspicious InfoPaths and highlights the Top-5 for further inspection, allowing analysts to focus on the most critical alerts and enhancing situational awareness prior to full investigation. In addition, the model conducts an in-depth analysis of detected command-line activity, offering clear explanations of intent and potential impact. This contextual intelligence supports more accurate threat identification, reduces false positives, and streamlines decision-making for security operations teams.
Figure 10: GPT-Powered Snapshot Alarm and Threat Prioritization: An overview of the GPT model’s recommendation process, which ranks suspicious InfoPaths and provides in-depth analysis of command-line activities.
V Evaluation
@@
V-A Attack Datasets
V-A1 DARPA Engagement Series Datasets (Linux and Unix-Like Servers)
@@
TABLE III: GitHub Availability, Reproducibility and Real-Time Detection of Related Work
@@
TABLE IV: Licensed Commercial EDR Solutions
@@
TABLE V: Node-Level Results on DARPA E3 Datasets
@@
TABLE VI: Attack Scenario Checklist (E3)
@@
V-A2 Industrial Real-Time Detection (Microsoft Windows End-User and Servers)

### v2

@@ III-A Attack-Clause Sketch: Graph Construction and Quantification
Figure 6: Interphase attack association: The Leiden community detection algorithm groups related process nodes, overcoming fragmentation issues to correlate multi-stage attacks.
III-A1 Refined Provenance Structure and Reduction
Traditional approaches often construct heterogeneous graphs where processes, files, sockets, and command strings are distinct nodes (Figure 4(a)). This results in high-dimensional, noisy graphs where semantic context is diluted.
To address this, we define a Refined Command-Line Focused Structure (Figure 4(b)). Let the system execution be represented as a directed graph $G=(V,E,\mathcal{A})$. Here, $V$ represents the set of uniform system process nodes. $E\subseteq V\times V$ denotes direct causal dependencies (e.g., process forking). Crucially, command-line arguments and network contexts are not separate nodes but are encapsulated as an attribute set $\mathcal{A}_{v}=\{\text{cmd}_{v},\text{net}_{v},\dots\}$ resident within each process node $v\in V$. This isomorphic structure preserves semantic richness while significantly reducing graph complexity for downstream embedding and traversal algorithms.
To mitigate the dependency explosion problem [10], we apply a two-stage graph reduction function $\Phi:G\to G^{\prime}$.
First, Semantic Pruning removes nodes lacking command-line relevance. We define a subgraph $G_{sub}\subset G$ as irrelevant if for all nodes $v\in G_{sub}$, the command attribute $\text{cmd}_{v}=\emptyset$. These subgraphs are pruned as they typically represent benign background noise.
Second, Acyclic Transformation ensures valid causal tracing by eliminating cyclic dependencies. $G^{\prime}$ is transformed into a Directed Acyclic Graph (DAG) via depth-first search based cycle removal, ensuring that any traversal represents a valid temporal flow of events.
III-A2 Causal Inference via Hybrid Node Scoring
Identifying critical nodes within massive provenance graphs requires quantifying their structural importance. We employ a hybrid scoring mechanism combining probabilistic ranking and centrality measures to determine initial node weights, as depicted in Figure 5.
Quantifying Rareness (Inverted PageRank). We adapt the PageRank algorithm [2] to identify anomalous nodes based on topological scarcity. Let $PR(v)$ denote the PageRank score of node $v$. The iterative computation is defined as:
$PR(v)_{t+1}=(1-d)+d\sum_{u\in\text{In}(v)}\frac{PR(u)_{t}}{L(u)}$ (1)
where $d\approx 0.85$ is the damping factor, $\text{In}(v)$ are incoming predecessors, and $L(u)$ is the out-degree of $u$. Unlike standard applications where high PR indicates prominence, in attack provenance, nodes with low in-degree (lower PR) often signify rare, potentially malicious execution paths. We normalize these scores to emphasize rarity.
Identifying Root Causes (Betweenness Centrality). To find pivotal ”bridge” nodes that connect disparate attack phases (e.g., initial access leading to privilege escalation), we utilize Shortest-Path Betweenness Centrality ($CB$) [3]. For a node $v$, $CB(v)$ quantifies the fraction of shortest paths between all node pairs $(s,t)$ that pass through $v$:
$CB(v)=\sum_{s\neq v\neq t}\frac{\sigma_{st}(v)}{\sigma_{st}}$ (2)
where $\sigma_{st}$ is the total number of shortest paths from $s$ to $t$, and $\sigma_{st}(v)$ is the count of those paths passing through $v$. High $CB$ scores highlight structurally critical bottlenecks.
Edge Weight Initialization. To facilitate path-based analysis, these node-level metrics must be projected onto the edges. For any directed edge $e_{mn}\in E^{\prime}$ connecting parent $m$ to child $n$, we define the initial Base Edge Weight ($W_{mn}^{base}$) as the normalized aggregate of the structural significance of both connected nodes:
$W_{mn}^{base}=\text{Norm}\left(PR(m)+PR(n)+CB(m)+CB(n)\right)$ (3)
where $\text{Norm}(\cdot)$ scales the resulting magnitude to the range $[0,1]$. This weight represents the inherent structural interest of the transition from $m$ to $n$.
III-A3 Interphase Attack Association via Community Detection
Attacks often manifest as distinct phases separated by time, making causal linkage difficult. To correlate these multi-stage behaviors, we partition the reduced graph $G^{\prime}$ into semantically coherent clusters.
We utilize the Leiden algorithm [16], an improvement over Louvain that guarantees connected communities and optimizes modularity. As illustrated in Figure 6, Leiden partitions the graph into a set of disjoint communities $\mathcal{C}=\{C_{1},C_{2},\dots,C_{k}\}$. Each detected community $C_{k}$ is assigned a Community Score ($S_{C_{k}}$) based on the aggregate risk density within that cluster. This score serves as a coherence multiplier in subsequent analysis, reinforcing links between activities occurring within the same attack phase context.
III-B Attack-Evidence Awareness: Detection and Reporting
The second phase leverages the weighted provenance graph to detect threats. This involves differentiating command-line semantics against known signatures and employing ensemble learning for unknown anomaly detection, as outlined in Figure 7.
Figure 7: Explainable anomaly detection mechanism: The process involves retrieving ranked InfoPaths, analyzing command-line differences via SimHash, and using a multi-classifier ”Police-Witness” voting policy for final verdict.
III-B1 Differentiation of Known Threats (Rule-Based Boosting)
To integrate expert knowledge, we evaluate process command-lines against curated threat intelligence, specifically Sigma rules [29] and defense evasion patterns [12]. Let $R(e_{mn})$ be a boolean indicator function that returns 1 if the command-line context $\text{cmd}_{n}$ associated with edge $e_{mn}$ matches a known threat signature, and 0 otherwise. We introduce a Risk Coefficient $\alpha>1$ corresponding to the severity of the matched rule. The base edge weight is updated to a Refined Edge Weight ($W_{mn}^{refined}$) to reflect this deterministic knowledge:
$W_{mn}^{refined}=\begin{cases}W_{mn}^{base}\times\alpha&\text{if }R(e_{mn})=1\\ W_{mn}^{base}&\text{otherwise}\end{cases}$ (4)
This operation significantly amplifies the weights of paths containing known malicious indicators, ensuring their prioritization during retrieval.
III-B2 Differentiation of Unknown Anomalies (SimHash Ensemble)
Detecting novel attacks requires identifying subtle deviations in command-line semantics, often disguised through obfuscation. We propose an approach combining semantic embedding with locality-sensitive hashing.
Semantic Hashing via SimHash. Standard embeddings (e.g., Word2Vec) result in high-dimensional vectors where similarity search is computationally expensive. To efficiently detect obfuscated variations, we apply SimHash [27]. Let $\mathbf{v}_{c}\in\mathbb{R}^{d}$ be the dense vector representation of a command string $c$. The SimHash function $H:\mathbb{R}^{d}\to\{0,1\}^{f}$ maps this vector to an $f$-bit binary signature $h_{c}$. A crucial property of SimHash is that the Hamming distance between two signatures approximates the cosine similarity of their original vectors:
$\text{Hamming}(h_{c_{i}},h_{c_{j}})\approx\theta(\mathbf{v}_{c_{i}},\mathbf{v}_{c_{j}})$ (5)
This allows us to efficiently quantify semantic deviation, as shown in Figure 8, where obfuscated commands yield small but detectable Hamming distances from benign counterparts.
Figure 8: Example of SimHashed command-line executions. SimHash maps semantically similar commands to nearby binary signatures, allowing detection of obfuscated variations (red box) via Hamming distance.
Ensemble ”Police-Witness” Voting. To avoid algorithmic bias inherent in single-model anomaly detection, we implement an ensemble voting strategy [25]. We employ a diverse set of unsupervised classifiers $\mathcal{M}=\{\text{iForest},\text{LOF},\text{OC-SVM},\text{EllipticEnvelope},\dots\}$. Each classifier $M_{j}\in\mathcal{M}$ evaluates the SimHashed InfoPaths independently. The final classification of a path $P$ is determined by a majority vote:
$V(P)=\mathbb{I}\left(\sum_{M_{j}\in\mathcal{M}}\mathbb{I}(M_{j}(P)=\text{anomalous})>\frac{|\mathcal{M}|}{2}\right)$ (6)
Paths confirmed as anomalous by the majority are flagged for analyst review.
III-B3 InfoPath Retrieval and GPT-Powered Reporting
The final step is to extract the most salient attack chains, termed InfoPaths, from the graph. An InfoPath is defined as a directed sequence of causally linked process nodes $(v_{1},v_{2},\dots,v_{k})$ representing a potential attack progression.
To retrieve the most critical paths, we adapt Dijkstra’s shortest-path algorithm [13]. Since standard Dijkstra minimizes cost, and our refined weights $W_{mn}^{refined}$ represent *evidence* (where higher is more significant), we define the traversal cost function as the inverse of the evidence, further weighted by the community score to favor cohesive attack phases:
$Cost(e_{mn})=\frac{1}{W_{mn}^{refined}\times S_{C_{community(n)}}}$ (7)
By minimizing this cost, the algorithm effectively maximizes the cumulative evidence along the path, prioritizing InfoPaths with high structural rarity, centrality, known threat indicators, and community coherence.
To mitigate alert fatigue, the Top-$k$ retrieved InfoPaths are fed into a Retrieval-Augmented Generation (RAG) framework [26] powered by Llama-2. The model augments the raw path data with context from external threat intelligence reports, generating a concise, prioritized snapshot alarm that explains the attack intent and recommended actions.
IV Evaluation
@@
IV-A Attack Datasets
IV-A1 DARPA Engagement Series Datasets (Linux and Unix-Like Servers)
@@
TABLE II: GitHub Availability, Reproducibility and Real-Time Detection of Related Work
@@
TABLE III: Licensed Commercial EDR Solutions
@@
TABLE IV: Node-Level Results on DARPA E3 Datasets
@@
TABLE V: Attack Scenario Checklist (E3)
@@
IV-A2 Industrial Real-Time Detection (Microsoft Windows End-User and Servers)